This week we had lots of visitors from Austin. We had new equipment to install on the tracker, for example, the field calibration unit. This is the device at the entrance aperture of the corrector which can shine light from flat lamps or from line lamps for calibration of our spectrographs.
We also had delivery of the LRS2-R spectrograph and a temporary installation of a VIRUS unit. The VIRUS unit was put in place mostly to test the mechanics of the enclosure cooling and control systems. We were able to get our first spectrum with both spectrographs. By the end of the week we were able to find a guide star with guide probes at one specific orientation and then hold a science target on a specific location of the LRS2-R and VIRUS IFUs. We then could move the target around the IFU to confirm the direction within the IFU. Even the weather cooperated finally, with some near photometric conditions we put a spectrophotometric standard on the LRS2-R. On the VIRUS unit we confirmed that the dither mechanism does work and we were able to move a target from one fiber to the adjacent fiber.
We will be moving back to guide probe commissioning in the coming week and may not get back to LRS2 instrument commissioning until January.
All in all a very exhausting, but good week!!!
This past week we had a large fraction of the Austin engineering and software groups visiting the telescope. This gave us an opportunity to make rapid improvements in the Telescope Control System and start getting the telescope ready for the installation of the first spectrograph.
With the removal of the Deployable Wave Front Sensors we were able to re-enable the guide probes and continue with their commissioning. Despite rather poor weather much of the week, the big milestone reached was the ability to now guide on the guide probes.
The software and engineering teams remain at the HET this next week and if we have some clear skies we will continue the push on commissioning of both the guide probes and refinement and characterization of the mount model.
We have big news this week. We have officially declared first light on the wide field upgrade for the HET. Please see:
Upgrade HET Sees First Light
This is the first light image from the 29th of July. The image quality is 1.3″ and we were only an arcminute from the target of interest.
Now that verification of the WFU corrector optics is complete we can get started on operational aspects of the new HET. This next week we are collecting data with the guide probes which will allow us to start the software commissioning of the guide probes next week.
We have cleared out the temporary clean room we installed in our loading bay so we can now have access from the mirror lab to the HET dome floor. This means that the mirror team will start doing mirror re-coating in the near future.
Other interesting news comes from our big hail storm in the past month. It broke many car windows damaged a few of the residents’ roofs and below we show the damage to the HET dome. Thankfully no leaks that we can detect.
In the big hail storm of 2015 that fully covered the road in several inches of hail the HET dome took a bit of a beating.
Since the last blog posting we had our most senior telescope operator retire. We wish her well in all of her next adventures. In the mean time we have posted the position and it is open now, probably for two more weeks.
Telescope Operator Position now Open